# An Algorithmic Approach to Finding Degree-Doubling Nodes in Oriented Graphs

### v6

@@ A Graph Level Order (GLOVER) on a set $S$ can be defined as follows:
To further illustrate the Graph Level Order (GLOVER) data structure, imagine a social media example where users can follow one another but no-one follows back. This network is then an oriented graph. Running the Seymour Second Neighborhood Conjecture in this context means looking for an account whose indirect sphere of influence is greater than their direct sphere of influence. Here, we are defining the indirect sphere of influence as those accounts they can reach through others, and we are defining the direct sphere of influence as those accounts they can directly connect with. Essentially, these are the same definitions as the second neighbor and first neighbor respectively that we have been working with. With this in mind, after finding the person who connects to the fewest accounts in this social media graph, we could set up containers just as before.
Now, lemma 5.4 (Interior Degree at least 1) says that each person in container 1 must follow another person in container 1. What this is saying is that, inside this first container, we will have one cycle of users where the messages are traversed about, or a number of disjoint cycles where messages are spread among themselves. Considering example 5.6, which takes the former of these approaches. These cycles in the first container are what stops the user Andi from being an influencer.
This brings us to the next container. Lemma 5.6 (Weak Decreasing Exteriors Lemma) lets us know that this container will have 5 people - (Hugo, Isla, Jade, Kai, and Lei). If any person in this container followed less than two other people in this same container, then they would need to follow at least five people in the next container (container 3). From this we would conclude that a person in container 1, whose indirect sphere of influence is this persons sphere of influence, would have its first degree double and its decond degree double, implying that its overall degree doubles.
@@
$\Rightarrow$ BellChadDawnElseFayeGus $\Rightarrow$ HugoIslaJadeKaiLei $\Rightarrow$ H
Figure 7: This is an example of a social media setup of the Seymour Second Neighborhood Conjecture. The Second Container has ten arcs and 5 nodes, which represents a perfect graph. This implies that in the next container $C_{3}$ (not shown), with 4 people can only have 6 arcs. Due to this fact, at least one person in $C_{3}$ will relate to at least five people in $C_{4}$. We can conclude then that at least one person in $C_{1}$ will be an influencer and have their degree double.
Table 1: Intersection Results for Triplets of Nodes in 5.6
Triplet(A, B, C) int(A, B) ext(B, C) Intersection Result
[…]
@@ 6 Back Arcs
With a theoretical foundation underway, we are ready to launch an investigation into back arcs. We will look into whether these arcs can exist in a counterexample and what other properties of oriented graphs must be true.
It is important to restate that back arcs and exterior neighbors are distinct concepts. Back arcs are not a subset of exterior neighbors. While back arcs connect to any previous container, exterior neighbors are limited to connecting with the next container. Additionally, exterior neighbors do not represent all possible external neighbors. Rather, an exterior neighbor ext(x, y) specifically refers to nodes at a distance of two from $x$ that are also neighbors of $y$.
@@
One of the things we notice about Lemma 4 (Minimum Out-Degree 3 with Neighbors 1) when $v_{0}$’s out-degree did not double in $G^{2}$, compared to the previous examples when it did, is that $v_{0}$ in Lemma 4 allowed for a cycle to exist in the out-neighbors of $v_{0}$. In such a situation, all $x\in N^{+}(v_{0})$ can still have $d^{+}(x)\geq d^{+}(v_{0})$ without the out-neighbors of $x$ causing $v_{0}$’s out-degree to double. This next lemma formalizes that concept.
@@
Assume that $u$ and $v$ are two nodes in container $C_{i}$, and that they share a parent $v_{p}$ in container $C_{i-1}$. Suppose that $|\text{int}(v_{p},v)|\geq|\text{int}(v_{p},u)|+1$
In order to preserve the necessary degree limits in the graph, $v$ will need fewer external connections if its inner degree is greater than $u$. $u$, on the other hand, will need more external connections to make up for its lack of inner connections.
As a result of the imbalance in internal degrees, the external connections of $u$ are essentially raising the second neighbors of degree of $v_{p}$.
As a result, the exterior degree of $v_{p}$ will double to make room for the additional connections that $u$ in the graph requires.
Thus, these different interior degrees of $u$ and $v$ in container $C_{i}$, will cause the exterior degree of $v_{p}$ in $C_{i-1}$ to double. This leads to a contradiction, so the claim holds. ∎
Lemma 6.1 (Regular Interior Degrees) provides another insight into the structure within containers of these counterexamples. In particular, it shows that all the out-neighbors of interior arcs within a container must have the same degree.
In Example 5.6 and Table 1, we examined a minimum degree node $v_{0}$ with degree 6 in a regular graph. The interior degrees of nodes increase outwardly: $v_{0}$ has an interior degree of 0, nodes in $C_{1}$ have 1, nodes in $C_{2}$ have 2, and so forth.
The definition of a back arc is an edge $(y,z)$, where $y\in C_{i}$, $z\in C_{j}$, and $j<i$ respectively. It is implied that the source node $y$ acquires new neighbors at a lower level, maybe doubling its degree, if such a back arc were included. The shortest-path lengths from $v_{0}$ strictly define each level in the hierarchical container organization, which is in contrast to this.
@@
Since $N^{+}(v_{k})$ includes all direct neighbors of $v_{k}$, it follows that:
$|N^{+}(v_{k})|=deg(v_{k}).$
By the structure of GLOVER and the absence of back arcs from $u_{i}$ to $v_{0},\dots,v_{k-1}$, the size of $ext(u_{i},v_{k})$ satisfies:
@@
Each node in $u_{i}$’s exterior is matched up with a node in $v_{k}$’s neighborhood. This means that the back arc will double $u_{i}$’s exterior degree. Then, because we’ve already assumed that $u_{i}$’s interior degree will double, we can say that $u_{i}$’s overall degree will double. This is true because its interior degree and its exterior degree will double.
@@
Back Arcs Special Case
Let $G$ be an oriented graph represented by a GLOVER, with containers $C_{1},C_{2},\ldots,C_{n}$ defined based on the distance from the minimum degree node $v_{0}$ in order to prevent a degree doubling node.
@@
First we must notice that the order within the container system, which defines exterior arcs as going to the next container (i.e. from container $C_{i}$ to container $C_{i+1}$) and not two containers down (i.e. not from container $C_{i}$ to container $C_{i+2}$), explicitly forbids $x$ having an exterior neighbor $z$ in the container $C_{i+2}$ where $x$ is in container $C_{i}$. This gives the only possibility for $z$ as being an interior neighbor of $x$.
@@
First Back Arc Between Consecutive Containers in an Oriented Graph:
Let $G$ be an oriented graph with minimum degree node $v_{0}$, with minimum degree $d$. Suppose that we represent this graph by a GLOVER. Suppose we are attempting to prevent a degree-doubling node and that all arcs point forward (i.e., from container $C_{j}$ to container $C_{j+1}$ for $j<i$) with no back arcs between containers $C_{1},C_{2},\dots,C_{i-1}$. If the first back arc occurs from container $C_{i}$ to $C_{i-1}$, then it will introduce an arc violating $G$’s oriented property.
@@
Now, let the first back arc appear from a node $u_{i}\in C_{i}$ to a node $v_{i-1}\in C_{i-1}$. In this oriented graph, by definition, all arcs are supposed to point in the forward direction—from container $C_{j}$ to container $C_{j+1}$. By Lemma 6.1 (Regular Interior Degrees), we know that all nodes inside container $C_{i-1}$ have the same interior degree. Because of their matching interior degrees, they will also share the same exterior degrees. Then since the node $v_{i-1}$ is in $C_{i-i}$, $v_{i-1}$ also maps to these exterior nodes. If $u_{i}$ is not mapped to by any node in $C_{i-1}$, then $dist{v_{0},u_{i}}-i$ so $u_{i}$ cannot belong to container $C_{i}$.
Therefore, the existence of a back arc from $C_{i}$ to $C_{i-1}$ contradicts the oriented nature of $G$, and no back arcs can exist prior to $C_{i}$.
Lemma 6.4 (No Back Arc to Early Container) is a special case where we can apply Lemma 6.1 (Regular Interior Degrees). From this, we can see that a back arc to a parent container violates the orientated nature of the graph.
@@
No Back Arcs Non-Existence of Back Arc
In an oriented graph $G$, represented by a GLOVER in an attempt to prevent a degree-doubling node, with a minimum degree node $v_{0}$ and containers $C_{1}$, $C_{2}$, $\dots$, and $C_{k}$, there cannot exist back arcs between a node $u_{i}\in C_{i}$ and a node $v_{j}\in C_{j}$, where $j<i$.
@@
This was proved in the lemmas 6.2 (No Back Arc to Early Container) and 6.4 (First Back Arc Between Consecutive Containers).
@@
Figure 9: Illustration of the proof of 7.2 (Interior Degrees Double). This is another illustration of the Interior Degree Doubling. Each node inside a container represents a node whose degree is doubled by the inductive hypothesis. This process cannot continue forever so we must reach the last node, and the parent of this last node in the sequence is the node whose exterior degree will double.
@@
[…]
We stated that back arcs could be detrimental to our algorithm, so the removal of them in this section was important. Now we can better navigate the other properties of oriented graphs.
7 Interior Degrees Doubling
This section introduces the concept of interior degree-doubling. This is similar to solving independent subproblems, where the subproblems we are attempting to solve here are the doubling of a node’s interior degree inside of a container. This can be seen as a localized version of Seymour’s Second Neighborhood Conjecture, but for individual containers instead of the overall graph.
@@ 7 Interior Degrees Doubling
Indistinguishable Regular Interior Nodes
Suppose we have a regular container $C$ of size $n$ of an oriented graph $G$ represented by a GLOVER. We are attempting to prevent a degree-doubling node. If a node $v_{p}$ ’s interior degree doubles, then every other node $v_{i}$ in the same container as $v_{p}$ will also have its interior degree doubled.
@@
Let $C_{i}$ be a regular container of size $n$. Let $v_{p}$ be a node in this container, where $int(u,v_{p})$ doubles in $C_{i-1}$ for some $u\in C_{i-1}$. The interior degree of a node $v_{i}$ in a container is defined as the number of arcs shared by $v_{i}$ between $v_{i}$ and a neighbor or parent node. In a regular container, all nodes are indistinguishable in terms of their degree. They also have the same parents because of the DSP. Any automorphism we construct, including a random automorphism $\phi$ of the container $C_{i}$, will preserve the degree structures. This means any node will have its interior degree doubled.
The crucial takeaway from Lemma 7.1 (Indistinguishable Regular Interior Nodes) is that regularity allows us to treat any node within a regular container as indistinguishable from others. Previously we were searching for a node in a container whose interior degree had doubled. That would be the key. What indistinguishable tells us is that if one node’s interior degree doubles inside a container, then all the node’s interior degrees double. That means that we can select any of those nodes.
Lemma 7.2.
Interior Degrees Double
Let $G$ be an oriented graph represented by a GLOVER. We are attempting to prevent a degree-doubling node. Let $v_{0}$ be a minimum-degree node in $G$ where all nodes satisfy the decreasing sequence property (DSP). Then, at least one node in each container $C_{i}$ will have its interior degree doubled.
Proof.
We prove this by strong induction on the size of the graph.
Base Case: In Lemmas 4.1 (Minimum Out-Degree ¡ 3), 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0), and 4 (Minimum Out-Degree 3 with Neighbors 1), we establish that for graphs with minimum degrees of 0, 1, 2, and 3, there exists at least one node in each container with a doubled interior degree. These cases form our base, supporting the induction.
@@
, where $u_{i}$ is the minimum degree node in $C_{i}$.
@@
By Lemma 5.6 (Weak Decreasing Exteriors Lemma), we know that $|H_{i}|>|H_{i+1}|$ for all $i>0$ resulting in a strictly decreasing sequence of positive integers. An increase in $int(x,y)$ requires a proportional reduction in $ext(y,z)$. This construction yields a strictly decreasing sequence of subgraph sizes: $|H_{1}|>...>|H_{k}|$.
The strictly decreasing nature of this sequence implies that it must terminate when $H_{i}\leq 2$. At this point, there are not enough arcs (1) in $H_{i}$ for both nodes to have an interior degree of 2. So, one node must have all its degree be exterior.
Consequently, for some node $y^{*}\in H_{i}$, we find that $int(x,y^{*})=0$ for some $x\in H_{i-1}$, leading to $ext(x,y^{*})=d^{+}(y^{*})$. Thus, the parent of $y^{*}$, denoted $y^{*}-1$, in $G$ achieves a doubled degree within $C_{i}$.
Therefore, for a graph of size $n$, at least one node in each $C_{i}$ attains a doubled interior degree, completing the inductive step.
Lemma 7.2 (Interior Degree Double) is a local version of Seymour’s Second Neighborhood Conjecture. We were able to isolate the containers as an oriented graph and searched for a node whose (interior) degree doubles. By finding a positive answer to this, and not just a positive answer, but a constructive algorithm that produces a positive answer, it gives us promise towards progress resolving the overall conjecture.
The way the proof was iterated was through both strong induction and a monotonically decreasing sequence of subsets. This cannot continue indefinitely. Once again, we see why the Decreasing Sequence Property (DSP) received its name.
This double of interior degrees will help us to build our algorithm in the search for a counterexample. It does, however, give us a bit of light into the side of the Conjecture being true.

### v7

@@ A Graph Level Order (GLOVER) on a set $S$ can be defined as follows:
Figure 7: Illustration of the Weak Decreasing Exteriors Lemma. We have a node $x$ in container $C_{i}$ connecting to three nodes in container $C_{i+1}$. In order to prevent $x$’s degree from doubling, we need to have $|C_{i+2}|<|C_{i+1}|$.
To further illustrate the Graph Level Order (GLOVER) data structure, imagine a social media example where users can follow one another, but no one follows back. This network is then an oriented graph. Running the SSNC in this context means looking for an account whose indirect sphere of influence is at least as large as their direct sphere of influence. Here, we are defining the indirect sphere of influence as those accounts they can reach through others. Similarly, we are defining the direct sphere of influence as those accounts they can directly connect with. Essentially, these are the same definitions as the second neighbor and first neighbor, respectively, that we have been working with. After finding the person who connects to the fewest accounts in this social media graph, we could set up containers just as before.
Now, lemma 5.4 (Interior Degree at least 1) says that each person in container 1 must follow another person in container 1. What this is saying is that, inside this first container, we will have one cycle of users where the messages are traversed about or a number of disjoint cycles where messages are spread among themselves. Considering example 5.5, which takes the former of these approaches. These cycles in the first container are what stops the user Andi from being an influencer.
This brings us to the next container. Lemma 5.6 (Weak Decreasing Exteriors Lemma) lets us know that this container will have 5 people—Hugo, Isla, Jade, Kai, and Lei. If any person in this container followed less than two other people in this same container, then they would need to follow at least five people in the next container (container 3). From this we would conclude that a person in container 1, whose indirect sphere of influence is this person’s sphere of influence, would have its first degree double and its second degree double, implying that its overall degree doubles.
@@
$\Rightarrow$ BellChadDawnElseFayeGusHugoIslaJade $\Rightarrow$ KaiLeiMiaNovaOwenPatQinRyan $\Rightarrow$ SageTateUgoVeraWadeXiaYoel $\Rightarrow$ H
Figure 8: This is an example of a social media setup of the SSNC. The second container has 21 arcs and seven nodes, which represents an orientation of the complete graph.
@@ 6 Back Arcs
With our discussion of the data structure concluded, we are ready to launch an investigation into back arcs. We will investigate whether these arcs can bring about a counterexample.
It is important to note that the back arcs and the exterior neighbors are distinct concepts. Back arcs are not a subset of exterior neighbors. While back arcs may connect to any previous container, exterior neighbors are limited to connecting with the next container. In addition, exterior neighbors do not represent all possible future containers. Rather, an exterior neighbor $ext(x,y)$ specifically refers to nodes at a distance of two from $x$ that are also distance one from $y$.
@@
One of the things we notice about Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1) when $v_{0}$’s out-degree did not double in $G^{2}$, compared to the previous examples when it did, is that $v_{0}$ in Lemma 4.3 allowed for a cycle to exist in the out-neighbors of $v_{0}$. In such a situation, all $x\in N^{+}(v_{0})$ can still have $d^{+}(x)\geq d^{+}(v_{0})$ without the out-neighbors of $x$ causing $v_{0}$’s out-degree to double. The next lemma formalizes that concept.
@@
Assume that $u$ and $v$ are two nodes in container $C_{i}$. Suppose that they share a parent $v_{p}$ in container $C_{i-1}$. Suppose also that $|\text{int}(v_{p},v)|\geq|\text{int}(v_{p},u)|+1$.
If $|int(v_{p},v)|>|int(v_{p},u)|+1$, then $v$ will have fewer exterior neighbors than $u$. Similarly, $u$ will have more external neighbors because of its fewer interior neighbors.
These exterior neighbors of $u$ are raising the second neighbors of $v_{p}$.
As a result, the exterior degree of $v_{p}$ will double. Because $v_{p}$ has already had its interior degree doubled, it will have its overall degree double
Thus, the different interior degrees of $u$ and $v$ in container $C_{i}$ leads to a contradiction, so the claim holds. ∎
Lemma 6.1 (Regular Interior Degrees) provides another insight into the structure within the containers of these counterexamples. It shows that all the out-neighbors of interior arcs within a container must have the same degree. This may not be surprising, as the examples we have been showing have all had this property. For example, both containers 1 and 2 in example 5.5 are orientations of regular graphs. The second container is actually an orientation of $K_{5}$, the complete graph on 5 nodes. This result begins to lend itself to some of the mathematical symmetry and structure that must be present in these containers and the SSNC.
The definition of a back arc is an edge $(y,z)$, where $y\in C_{i}$, $z\in C_{j}$, and $j<i$ respectively. It is implied that the source node $y$ acquires new neighbors at a lower level, maybe doubling its degree, if such a back arc were included. The shortest-path lengths from $v_{0}$ strictly define each level in the hierarchical container organization, which contrasts with this.
Next, we begin to look at how back arcs interact with the data structure and the development of the algorithm. Up to this point, we have gotten around back arcs by assuming that they do not exist. However, that will not last forever. These next few lemmas address the possibility of what happens when we try to construct a GLOVER and encounter a back arc. We will see that although we are not able to construct a GLOVER, we are still able to execute our algorithm and find a degree-doubling node.
Consider the following table based on example 5.5.
Table 1: Consequences of Back Arcs in 5.5
Back Arc From(A) Back Arc To (B) Back Arc New Neighbors via $N^{+}(B)$
Sage Faye Gus, Kai, Lei, Mia, Nova, Owen, Pat, Qin, Ryan
Tate Gus Hugo, Kai, Lei, Mia, Nova, Owen, Pat, Qin, Ryan
Ugo Hugo Isla, Kai, Lei, Mia, Nova, Owen, Pat, Qin, Ryan
Vera Bell Chad, Kai, Lei, Mia, Nova, Owen, Pat, Qin, Ryan
Wade Chad Dawn, Kai, Lei, Mia, Nova, Owen, Pat, Qin, Ryan
Xia Dawn Else, Kai, Lei, Mia, Nova, Owen, Pat, Qin, Ryan
Yoel Else Faye, Kai, Lei, Mia, Nova, Owen, Pat, Qin, Ryan
This table looks at the possible back arcs from container 3 in example 5.5 and how they would impact the degree doubling of the node in question. The back arcs constructed in this table are attempts to go from container 3 $C_{3}$ to container $C_{1}$. In each case, we can see that $N^{+}(B)$ causes that degree of the source node to double. What follows is a lemma that generalizes this table.
@@
We know there are no back arcs from $u_{i}$ to $v_{0},\dots,v_{k-1}$, so the size of $ext(u_{i},v_{k})$ satisfies:
@@
This means that the back arc will double $u_{i}$’s exterior degree. Then, because we’ve already assumed that $u_{i}$’s interior degree will double, we can say that $u_{i}$’s overall degree will double. This is true because its interior degree and its exterior degree will double. ∎
@@
Back Arcs Special Case Let $G$ be an oriented graph represented by a GLOVER, with containers $C_{1},C_{2},\ldots,C_{n}$ defined based on the distance from the minimum degree node $v_{0}$ in order to prevent a degree doubling node.
@@
First we must notice that the order within the container system, which defines exterior arcs as going to the next container (i.e., from container $C_{i}$ to container $C_{i+1}$) and not two containers down (i.e., not from container $C_{i}$ to container $C_{i+2}$), explicitly forbids $x$ having an exterior neighbor $z$ in the container $C_{i+2}$ where $x$ is in container $C_{i}$. This gives the only possibility for $z$ as being an interior neighbor of $x$.
@@
First Back Arc Between Consecutive Containers in an Oriented Graph: Let $G$ be an oriented graph with minimum degree node $v_{0}$, with minimum degree $d$. Suppose that we represent this graph by a GLOVER. Suppose we are attempting to prevent a degree-doubling node and that all arcs point forward (i.e., from container $C_{j}$ to container $C_{j+1}$ for $j<i$) with no back arcs between containers $C_{1},C_{2},\dots,C_{i-1}$. If the first back arc occurs from container $C_{i}$ to $C_{i-1}$, then it will introduce an arc violating $G$’s oriented property.
@@
Now, let the first back arc appear from a node $u_{i}\in C_{i}$ to a node $v_{i-1}\in C_{i-1}$. In this oriented graph, by definition, all arcs are supposed to point in the forward direction—from container $C_{j}$ to container $C_{j+1}$. By Lemma 6.1 (Regular Interior Degrees), we know that all nodes inside container $C_{i-1}$ have the same interior degree. Because of their matching interior degrees, they will also share the same exterior degrees. Then since the node $v_{i-1}$ is in $C_{i-i}$, $v_{i-1}$ also maps to these exterior nodes. If $u_{i}$ is not mapped to by any node in $C_{i-1}$, then $dist(v_{0},u_{i})-i$ so $u_{i}$ cannot belong to container $C_{i}$.
Therefore, the existence of a back arc from $C_{i}$ to $C_{i-1}$ contradicts the oriented nature of $G$, and no back arcs can exist prior to $C_{i}$. ∎
Lemma 6.4 (No Back Arc to Early Container) is a special case where we can apply Lemma 6.1 (Regular Interior Degrees). From this, we can see that a back arc to a parent container violates the oriented nature of the graph.
@@
No Back Arcs Non-Existence of Back Arc In an oriented graph $G$, represented by a GLOVER in an attempt to prevent a degree-doubling node, with a minimum degree node $v_{0}$ and containers $C_{1}$, $C_{2}$, $\dots$, and $C_{k}$, there cannot exist back arcs between a node $u_{i}\in C_{i}$ and a node $v_{j}\in C_{j}$, where $j<i$.
@@
This was proved in the lemmas 6.2 (No Back Arc to Early Container) and 6.4 (First Back Arc Between Consecutive Containers). ∎
@@
Figure 10: Illustration of the proof of 7.2 (Interior Degrees Double). This is another illustration of the Interior Degree Doubling. Each node inside a container represents a node whose degree is doubled by the inductive hypothesis. This process cannot continue forever, so we must reach the last node, and the parent of this last node in the sequence is the node whose exterior degree will double.
@@
We stated that back arcs could be detrimental to our data structure. Now that we have shown that even if they are present, they will not hinder the development of our pathfinding algorithm, as a back arc will simply lead to a degree-doubling node faster. Now we can better navigate the other properties of oriented graphs.
7 Interior Degrees Doubling
This section introduces the concept of interior degree-doubling. This is like solving independent subproblems, where the subproblems we are attempting to solve here are the doubling of a node’s interior degree inside of a container. This can be seen as a localized version of Seymour’s Second Neighborhood Conjecture, but for individual containers instead of the overall graph.
Lemma 7.1.
Indistinguishable Regular Interior Nodes Suppose we have a regular container $C$ of size $n$ of an oriented graph $G$ represented by a GLOVER. We are attempting to prevent a degree-doubling node. If a node $v_{p}$ ’s interior degree doubles, then every other node $v_{i}$ in the same container as $v_{p}$ will also have its interior degree double.
Proof.
Let $C_{i}$ be a regular container. Let $v_{p}$ be a node in this container, where $int(u,v_{p})$ doubles in $C_{i-1}$ for some $u\in C_{i-1}$. In a regular container, all nodes are indistinguishable in terms of their degree. They also have the same parents because of containers. Any automorphism we construct, including a random automorphism $\phi$ of the container $C_{i}$, will preserve the degree structures. This means any node will have its interior degree doubled under the automorphism $\phi$. ∎
The crucial takeaway from Lemma 7.1 (Indistinguishable Regular Interior Nodes) is that regularity allows us to treat any node within a regular container as indistinguishable from others. Previously we were searching for a node in a container whose interior degree had doubled. That would be the key. What indistinguishable tells us is that if one node’s interior degree doubles inside a container, then all the nodes within that container will have their interior degrees double. That means that we can select any of those nodes.
We still need to prove that for each container there exists a node whose interior degree doubles. We have seen this in the example 5.5, but those could be too simple to observe any real patterns. In similar structure, what we saw in 4.1 (Minimum Out-Degree $<$ 3) was that it was easy to find degree-doubling nodes in graphs with a minimum degree of 0, 1, or 2. However, when that minimum degree was increased to 3, there were more cases that needed to be considered.
@@ 7 Interior Degrees Doubling
Interior Degrees Double Let $G$ be an oriented graph represented by a GLOVER. We are attempting to prevent a degree-doubling node. Let $v_{0}$ be a minimum-degree node in $G$ where all nodes satisfy the decreasing sequence property (DSP). Then, at least one node in each container $C_{i}$ will have its interior degree doubled.
@@
We prove this by strong induction on the size of the graph. Base Case: In Lemmas 4.1 (Minimum Out-Degree $<$ 3), 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0), and 4.3 (Minimum Out-Degree 3 with Neighbors 1), we establish that for graphs with minimum degrees of 0, 1, 2, and 3, there exists at least one node in each container with a doubled interior degree. These cases form our base, supporting the induction.
@@
where $u_{i}$ is the minimum degree node in $C_{i}$.
@@
By Lemma 5.6 (Weak Decreasing Exteriors Lemma), we know that $|H_{i}|>|H_{i+1}|$ for all $i>0$ resulting in a strictly decreasing sequence of positive integers.
This construction yields a strictly decreasing sequence of subgraph sizes: $|H_{1}|>...>|H_{k}|$.
The strictly decreasing nature of this sequence implies that it must terminate when $H_{i}\leq 2$. At this point, there are not enough arcs in $H_{i}$ for both nodes to have an interior degree of 2. So one node must have all of its degree be exterior.
Consequently, for some node $y^{*}\in H_{i}$, we find that $int(x,y^{*})=0$ for some $x\in H_{i-1}$, leading to $ext(x,y^{*})=d^{+}(y^{*})$. Thus, the parent of $y^{*}$, denoted $y^{*-1}$, in $G$ achieves a doubled degree in $C_{i}$.
Therefore, at least one node in each $C_{i}$ attains a doubled interior degree, completing the inductive step. ∎
Lemma 7.2 (Interior Degree Double) is a local version of Seymour’s Second Neighborhood Conjecture. We were able to isolate the containers as an oriented graph and searched for a node whose (interior) degree doubles. By finding a positive answer to this, and not just a positive answer but a constructive algorithm that produces a positive answer, it gives us promise towards progress resolving the overall conjecture.
The way the proof was iterated was through both strong induction and a monotonically decreasing sequence of subsets. This cannot continue indefinitely. This gives rise to why the Decreasing Sequence Property (DSP) received its name. Prior to proofs, our research showed the Weak Decreasing Exteriors Lemma as an early observation through many examples. The questions of whether this was always true or whether back arcs like lemma 6.3 could force this into being a false positive.
This proof also helps with explaining the concept of the SSNC. Any oriented graph in which we want to seek a degree-doubling node can be represented in a GLOVER, as long as there are no back arcs. Once we have done this, there are containers spread out from a chosen minimum degree node. Now for each container we can simply run the conjecture on that container. The difference is that with lemma 7.2 (Interior Degrees Double), we already know that there is a node whose degree doubles inside that container. Then with lemma 7.1 (Indistinguishable Regular Interior Nodes), we know that any node we select in that container will also have their interior degree double.
This notion of interior degrees will help us to build our algorithm in the search for a counterexample. It also gives us a bit of light into the possibility of the conjecture being true.